Find the slope and the y intercept of 8x=5y+35
Answer by lwsshak3(11628)

Find the slope and the y intercept of
8x=5y+35
convert to standard form of equation for a line: y=mx+b, m=slope, b=y-intercept
5y=8x-35
divide by 5
y=(8/5)x-7
slope=8/5
y-intercept=-7